OpenStudy (anonymous):

is it possible to solve: 7d^2+12=0??

OpenStudy (saifoo.khan):

Yes!

OpenStudy (saifoo.khan):

but the solutions will not be real.

OpenStudy (saifoo.khan):

\[d^2 = -\frac{12}{7} \]Now take square root on both sides.

OpenStudy (anonymous):

If we assume that \(d\) is a real number, then no. If it is allowed to be an imaginary or complex number, then most certainly!

OpenStudy (anonymous):

okay. i must've done the previous steps wrong. because the answer is -2 and 2

OpenStudy (anonymous):

umm can you solve 10d-2t=-40? then substitute it into 7d^2+5d-t-8=0 pleasse?

OpenStudy (saifoo.khan):

d and t?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

WolframAlpha is a good resource to check your answers: http://www.wolframalpha.com/input/?i=10d-2t%3D-40+and+7d%5E2%2B5d-t-8%3D0

OpenStudy (anonymous):

Check your work carefully... your troubles were likely caused by an erroneous slip of the pen. (It happens to all of us!)

OpenStudy (anonymous):

thanks, but it doesn't show how to solve it algebraically. which I need to do

jimthompson5910 (jim_thompson5910):

10d-2t=-40 -2t=-40-10d t = -40/(-2) - 10d/(-2) t = 20+5d <<--- You probably made the error here ------------------- 7d^2+5d-t-8=0 7d^2+5d-(20+5d)-8=0 7d^2+5d-20-5d-8=0 7d^2-28=0 Now solve for d

OpenStudy (anonymous):

ohhh ok. yeah i didnt know how to solve for d then I figured it out...thanks!

jimthompson5910 (jim_thompson5910):

you're welcome